# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ 3.2 Finetuning
Benchmark datasets. For iEEG downstream evaluation, we use Neuroprobe (2024neuroprobe), which provides 15 auditory, visual, and language decoding tasks from naturalistic movie-watching iEEG  and MAYO dataset (bbrinkm2014upenn)(seizure detection). For EEG evaluation, we use: FACED (chen2023faced) (emotion decoding), PhysioNet-MI (goldberger2000physionet) (motor-imagery classification), and MentalArithmetic (zyma2019electroencephalograms) (cognitive-workload decoding). Refer to Appendix F.2 for more detail.
Baseline models. We evaluate DIVER-1 against state-of-the-art foundation models across both modalities: LaBraM (jiang2024labram) and CBraMod (wang2024cbramod) for EEG, and BrainBERT (wang2023brainbert), Population Transformer (PopT) (chau2025popt) and Brant (zhang2023brant) for iEEG. These transformer-based models employ self-supervised pretraining strategies; LaBraM, CBraMod, BrainBERT and Brant use masked patch reconstruction objectives while PopT applies discriminative self-supervised learning on BrainBERT embeddings.
Finetuning method. For iEEG downstream tasks, we evaluate both linear probing and full finetuning using a linear classifier head on the flattened token representations. Linear probing freezes the encoder and trains only the classifier, providing a clean measure of representation quality. Full finetuning updates all encoder parameters together with the linear head. Details are in Appendix B.8. For EEG downstream tasks, we follow CBraMod’s finetuning protocol (wang2024cbramod), jointly training the encoder and a three-layer MLP classifier. This MLP generally outperforms a linear head on EEG benchmarks, and depth-dependent results are provided in Appendix Table 22.
@@ 4.1 Scaling Law
We systematically investigate scaling laws across multiple dimensions to understand how EFMs scale with computational resources and data. Our analysis evaluates performance on (1) pretext task loss (reconstruction loss during self-supervised pretraining) and (2) downstream task performance across diverse neural decoding benchmarks. We vary traditional scaling axes (compute budget, dataset size, model size) and EFM-specific factors (training epochs, subject diversity).
Overall scaling law validation. Both iEEG and EEG EFMs exhibit precise neural scaling behavior consistent with established Kaplan scaling laws (kaplan2020scaling), as shown in Figure 2 (a-c, e-g). We fit our results to the data-constrained scaling law framework (data_constrained_scaling_muennighoff2023scaling) (see Appendix Section A.2 for detailed scaling law background), providing the first quantitative evidence that scaling EFMs across compute, dataset size, model parameters, and training epochs – a dimension largely unexplored in EFM scaling – predictably and logarithmically improves performance. The scaling relationships exhibit strong log-log fits in iEEG EFMs, with R² values of 0.8152 for patch size 1 second models and 0.7718 for patch size 0.1 second models (Appendix Table 7), confirming the validity of the power-law scaling framework. The additional scaling result on EEG EFMs can be found at Appendix C.5.
Our empirical results follow the data-constrained scaling law (data_constrained_scaling_muennighoff2023scaling):
@@
where $N^{\prime}$ and $D^{\prime}$ account for diminishing returns with more epochs:
$D^{\prime}=U_{D}+U_{D}R_{D}^{*}\Big(1-e^{-\tfrac{R_{D}}{R_{D}^{*}}}\Big),\quad N^{\prime}=U_{N}+U_{N}R_{N}^{*}\Big(1-e^{-\tfrac{R_{N}}{R_{N}^{*}}}\Big)$ (8)
where $U_{D}$ , $R_{D}$, and $R_{D}^{*}$ each corresponds to unique data tokens, repetitions (epochs1-1), and the “half-life” of repeated data. The fitted parameters are detailed in Appendix C.3.
Standard scaling axes (model, data, compute). We observe in Figure 2 (a-c, e-g) expected power-law relationships across standard dimensions. Model size, dataset size, and compute budget scaling logarithmically reduces loss, following kaplan2020scaling, though Ephys data limitations necessitate data-constrained formulation.
EFM-specific scaling axes (epoch, subject number). Beyond the traditional dimensions of model size, dataset size, and compute budget, we identify two novel scaling dimensions relevant to Ephys.
@@
Epoch scaling in Figure 2 (d,h) shows increasing epochs improves performance across all model sizes, and larger models achieve lower loss when given sufficient repetitions. However, overly large models (e.g., XXL in Figure 2 (d) and Large in 2 (h)) perform worse at very small epoch counts, only surpassing smaller models after many more epochs—consistent with the data-constrained scaling law, where repeated passes increase the effective dataset size $D^{\prime}$ required for large-capacity models.
@@
Subject diversity scaling in Figure 2 (i) shows that under constant total training data volume, there exists an optimal balance between the number of subjects and data per subject.
@@
IsoLoss Analysis. While increases in model size and training epochs both improve performance, they also raise compute cost, making it necessary to balance the two. To study this trade-off under fixed compute, we use the empirical IsoLoss landscape with IsoCompute curves (Figure 2(p), left), which map parameter–epoch pairs to equal FLOPs. The contours reveal a clear trend: at any fixed compute level, smaller and mid-sized models achieve lower loss than larger models. Thus, within realistic compute budgets, allocating resources toward training smaller models longer is more effective than briefly training very large models. We revisit this in the Practical Implications section using predicted IsoLoss contours (Figure 2(p), right).
@@
Data Constrained Scaling Law Fitting. Our fitted data-constrained scaling law parameters reveal important domain-specific characteristics (Appendix Table 7). Most notably, EFMs exhibit smaller $R_{N}^{*}$ values (3.39 and 0.72 for 1s and 0.1s patch sizes, respectively) compared to language models (5.30) (data_constrained_scaling_muennighoff2023scaling). This indicates that increasing model parameters yields diminishing returns more rapidly in Ephys modeling than in language modeling, consistent with our finding that smaller models often suffice for EFM tasks.
@@
Downstream performance scaling. Following pretraining loss scaling, downstream performance (Figure 2 (i-o)) also scales across most dimensions. However, model size scaling exhibits different behavior across modalities: on iEEG tasks, larger models achieve similar performance to smaller models, while on EEG tasks, performance improves up to 813M parameters. The limited effect of model size on iEEG performance may reflect the simplicity of the binary classification tasks used in this benchmark, though comprehensive analysis are needed to fully characterize this behavior.
@@
Practical implications. These findings have direct consequences for resource allocation in EFM. Because the domain is data-constrained, the optimal compute strategy differs from that of language or vision: smaller models trained for more epochs outperform larger models trained briefly under fixed compute budgets. The IsoLoss and IsoCompute structure in Figure 2(p) (left) indicates that large models are compute-inefficient at realistic epoch counts. Building on this observation, Figure 2(p) (right) defines a clear “compute-optimal frontier” that specifies the most efficient combinations of model size and training duration. Prior EFMs tend to fall outside this frontier, often emphasizing increased model size over additional training repetitions (estimation details in Appendix G). Our scaling framework therefore provides a principled tool for selecting compute-optimal configurations before launching expensive pretraining runs.
It should be noted, however, that when the goal is to achieve the highest possible performance for a given model size, training beyond the compute-optimal frontier is still beneficial. This was the case for our Small iEEG model, and also appears to have been the case for CBraMod (wang2024cbramod) and BIOT (yang2023biot), which also trained substantially past their compute-optimal points.
Figure 2: Scaling laws and downstream performance of DIVER-1. (a-h) Scaling law validation: DIVER-1 follows data-constrained scaling laws across four dimensions for iEEG (a-d) and EEG (e-h) modalities. Loss decreases predictably with increased (a,e) compute (training FLOPs), (b,f) dataset size (number of tokens), (c,g) model size (parameters), and (d,h) training epochs, with strong log-log fits. iEEG experiments (a-d) used 100% of the dataset, while EEG experiments used 20% of the dataset for (e,h) and 100% for (f,g). (q-t) Downstream performance: Performance for iEEG (i-l) and EEG (m-o) across increasing (i) number of subjects while keeping dataset size identical, (j,m) dataset size (k,n) model sizes and (l,o) epochs. (p) Compute-Optimal Frontier (IsoLoss analysis): Comparison between empirical isolation loss contours and predicted isolation loss contours, with model configurations plotted to show the relationship between training epochs and model parameters under fixed compute budgets. (q) Neuroprobe benchmark results Comprehensive performance (AUROC) comparison across multiple neural decoding tasks, with $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ achieving state-of-the-art or competitive results on most tasks. $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ with $d_{\text{model}}=256$ and patch size 0.1s was pretrained on iEEG dataset for 32 epochs, past the compute optimal frontier for best performance. Performance with linear probing (red) and full finetuning (blue) are shown. (r, s) iEEG downstream performance (r) Neuroprobe multi-label classification results using $\textsc{DIVER}_{\text{Small}/\text{I}/0.1s}$. (s) MAYO(seizure detection task) results using $\textsc{DIVER}_{\text{Small}/\text{I}/1s}$ (t). EEG downstream performance: DIVER-1 showed competitive performance compared to other EEG foundation models (CBraMod and LaBraM-base) on the FACED, PhysioNet-MI, and MentalArithmetic datasets. Results shown are obtained using full finetuning. The DIVER model refers to $\textsc{DIVER}_{\text{Small}/\text{IE}/1s}$ with $d_{\text{model}}=512$ and patch size 1s pretrained on iEEG and EEG datasets for 16 epochs. Other baseline results are replicated using their official code. Performance values for CBraMod and LaBraM are reported from their original publications.
@@ 4.2 Decoding Performance
As shown in Figure 2(q,r,s,t), DIVER-1 iEEG and EEG each achieve state-of-the-art decoding performance across iEEG and EEG benchmarks. On the iEEG downstream dataset, our 13M-parameter model surpasses nearly all prior approaches—including BrainBERT (43M) (wang2023brainbert), PopT (63M)(chau2025popt), and Brant (506M) (zhang2023brant)—across nearly all 15 tasks in Neuroprobe binaray-label and by a large margin in MAYO. For Neuroprobe Multi-label, our model still surpasses the linear STFT laplacian, which surpasses other models except ours in binary-label. On EEG downstream tasks, DIVER-1 also establishes new SOTA results, outperforming CBraMod (wang2024cbramod) and LaBraM (jiang2024labram) on FACED emotion recognition (chen2023faced), PhysioNet-MI motor imagery (goldberger2000physionet), and MentalArithmetic workload decoding (zyma2019electroencephalograms). Full task-level values for all benchmark results are provided in Appendix Section D.1. We also evaluated DIVER-1 on additional EEG downstream tasks, where we identified a methodological (reproduction) problem. To ensure fair comparison and facilitate reproduction, we conducted a controlled investigation by standardizing the finetuning procedure across models and compring performance over seven different finetuning methods. We find that under controlled finetuning settings, DIVER-1 also achieves comparable or superior performance to the other baseline models (Appendix Table 24). Detailed experimental procedures and results are provided in Appendix Section E
The strong performance of NeuroProbe results is particularly notable given the pretraining–finetuning distribution shift under which DIVER-1 was trained. Unlike PopT or BrainBERT, DIVER-1 was pretrained on different datasets than those used for finetuning, on adult data rather than pediatrics, and using both ECoG and SEEG modalities rather than exclusively SEEG data. Moreover, DIVER-1 is much smaller (13M) than the other models (BrainBERT : 43M, PopT : 63M). This highlights the robustness of our representations across dataset shifts and recording modalities.
Interestingly, linear probing outperforms full fine-tuning on Neuroprobe. The limited sample size (-1750 for each fold) may be causing overfitting during full fine-tuning. Another possibility is the relatively low difficulty of the task. However, given that linear probing also outperforms in the multi-label, more challenging targets may be needed (e.g., regression to GPT-derived embeddings).
@@ 4.3 Ablation Studies
Because our performance gains reflect both architectural innovations and a much larger pretraining dataset, we disentangle these factors through architecture ablations and comparisons against baseline models trained on the same data.
Architecture ablations (Tables 17, 18) show that most components—any-variate attention, RoPE, STCPE, and the multi-domain reconstruction objective—generally improve performance across modalities. For iEEG specifically, several channel-wise embedding components (modality/subtype in some tasks and 3D position) produced slight improvements when removed, whereas in EEG these components typically boosted performance. This modality-dependent behavior suggests that future work may benefit from exploring architectures specialized for iEEG versus EEG.
To isolate architectural effects from data scale, we additionally compare models pretrained on the same dataset as existing baselines (Table 19). When pretrained on the BrainTreebank dataset used by BrainBERT and PopT, DIVER still achieves higher downstream performance, indicating that the architectural design itself yields stronger representations independent of data volume.
Further ablations and interpretability analysis are in Appendix D due to space constraints.

### v2

@@ 3.2 Finetuning
Benchmark datasets. For iEEG downstream evaluation, we use Neuroprobe (Zahorodnii et al., 2025), which provides 15 auditory, visual, and language decoding tasks from naturalistic movie-watching iEEG, and MAYO dataset (Bbrinkm and Cukierski, 2014)(seizure detection). For EEG, we use: FACED (Chen et al., 2023) (emotion decoding), PhysioNet-MI (Goldberger et al., 2000) (motor-imagery decoding), and MentalArithmetic (Zyma et al., 2019) (cognitive-workload decoding). Refer to Appendix F.2 for more details.
Baseline models. We evaluate DIVER-1 against state-of-the-art foundation models across both modalities: LaBraM (Jiang et al., 2024) and CBraMod (Wang et al., 2024c) for EEG, and BrainBERT (Wang et al., 2023), Population Transformer (PopT) (Chau et al., 2025) and Brant (Zhang et al., 2023) for iEEG. These transformer-based models employ self-supervised pretraining strategies; LaBraM, CBraMod, BrainBERT, and Brant use masked patch reconstruction objectives while PopT applies discriminative self-supervised learning on BrainBERT embeddings.
Finetuning methods. For iEEG downstream tasks, we evaluate both linear probing and full finetuning using a linear classifier head on the flattened token representations. Linear probing freezes the encoder and trains only the classifier, providing a clean measure of representation quality. Full finetuning updates all encoder parameters together with the linear head. Details are in Appendix B.8. For EEG downstream tasks, we follow CBraMod’s finetuning protocol (Wang et al., 2024c), jointly training the encoder and a three-layer MLP classifier. This MLP generally outperforms a linear head on EEG benchmarks, and depth-dependent results are provided in Appendix Table 23.
Figure 2: Contrasting scaling laws between DIVER-1 and LLMs. Each colored region shows how much each factor should be scaled (assuming compute optimal scaling) (multiplicative contribution) when total compute increases (with epoch fixed to two). Both DIVER-1 variants (left (iEEG) and middle (EEG)) demonstrates that data is the critical bottleneck for EFMs, requiring proportionally much larger increases in dataset size relative to model size. In contrast, LLM training (right) (adapted from Kaplan et al. (2020)) optimally allocates increased compute primarily to model size with minimal data growth, indicating that EFM are fundamentally more data-limited than LLMs.
@@ 4.1 Scaling Law
We systematically investigate multi-dimensional scaling laws to understand how EFMs scale with computational resources and data. Our analysis primarily uses (1) pretext-task loss (reconstruction loss during self-supervised pretraining) and (2) downstream-task performance across diverse Ephys benchmarks. We vary both traditional scaling axes (compute budget, dataset size, model size) and EFM-specific factors (training epochs, subject diversity).
Overall scaling law validation. Both iEEG and EEG EFMs exhibit neural scaling behaviors that aligns with the established Kaplan scaling laws (Kaplan et al., 2020), as shown in Figure 3 (a-c, e-g). Our empirical studies further show they also follow the more general data-constrained scaling law framework (Muennighoff et al., 2023):
@@
where $N^{\prime}$ and $D^{\prime}$ are diminishing returns with more epochs:
$\begin{split}D^{\prime}&=U_{D}+U_{D}R_{D}^{*}\Big(1-e^{-\tfrac{R_{D}}{R_{D}^{*}}}\Big),\\ N^{\prime}&=U_{N}+U_{N}R_{N}^{*}\Big(1-e^{-\tfrac{R_{N}}{R_{N}^{*}}}\Big).\end{split}$ (8)
where $U_{D}$ , $R_{D}$, and $R_{D}^{*}$ corresponds to unique data tokens, repetitions (epochs1-1), and the “half-life” of repeated data, respectively (see Appendix A.2 and C.3 for details). This set-up provides quantitative evidence that scaling EFMs across compute, dataset size, model parameters, and training epochs – a dimension largely unexplored in EFM scaling – predictably and logarithmically improves performance. The scaling relationships derived with DIVER-1 exhibit strong log-log fits in iEEG EFMs, with $R^{2}$ values of 0.8152 for patch size 1 second models and 0.7718 for patch size 0.1 second models (Appendix Table 8), confirming the validity of the power-law scaling framework (additional results in Appendix C.5). • Standard scaling axes (model, data, compute): We observe in Figure 3 (a-c, e-g) the expected power-law relationships across standard dimensions. Model size, dataset size, and compute budget scaling logarithmically reduces loss, following Kaplan et al. (2020), though Ephys data limitations necessitate data-constrained formulation. • EFM-specific scaling axes (epoch, subject number): Beyond the traditional dimensions of model size, dataset size, and compute budget, we identify two novel scaling dimensions relevant to Ephys.
Finding 1.
@@
Epoch scaling in Figure 3 (d,h) shows increasing epochs improves performance across all model sizes, and larger models achieve lower loss when given sufficient repetitions. However, overly large models (e.g., XXL in Figure 3 (d) and Large in Figure 3 (h)) perform worse at very small epoch counts, only surpassing smaller models after many more epochs—consistent with the data-constrained scaling law, where repeated passes increase the effective dataset size $D^{\prime}$ required for large-capacity models.
Finding 2.
@@
Subject diversity scaling in Figure 3 (i) shows that under fixed training data volume, there exists an optimal balance between the number of subjects and data per subject.
Finding 3.
@@
While increases in model size and training epochs both improve performance, they also raise compute level, making it necessary to balance the two. To study this trade-off under fixed compute, we use the empirical IsoLoss landscape with IsoCompute curves (Figure 3(p), left), which map parameter–epoch pairs to equal FLOPs. The contours reveal a clear trend: at any fixed compute level, smaller and mid-sized models achieve lower loss than larger models. Thus, within realistic compute budgets, allocating resources toward training smaller models longer is more effective than briefly training very large models. We revisit this in the Practical Implications section using predicted IsoLoss contours (Figure 3(p), right).
Finding 4.
@@
Our fitted data-constrained scaling law parameters reveal important domain-specific characteristics (Appendix Table 8). Most notably, EFMs exhibit smaller $R_{N}^{*}$ values (3.39 for 1s patch sizes and 0.72 for 1s) compared to language models (5.30) (Muennighoff et al., 2023). This indicates that increasing model parameters yields diminishing returns more rapidly in Ephys modeling than in language modeling, consistent with our observations that smaller models often suffice for EFM tasks.
Finding 5.
@@
Downstream performance.
Downstream performance scaling roughly mirrors pretraining behavior. Scaling with dataset size yields consistent gains, whereas scaling model size produces mixed effects (Figure 3(j,k,m,n)), similar to Anonymous (2026). In iEEG tasks, larger models perform comparably to smaller ones, while in EEG tasks performance continues to improve up to 813M parameters. Across training duration, performance generally improves up to 32 epochs for both iEEG and EEG tasks, although iEEG exhibits a slight decline at 64 epochs (Figure 3(i,o)). The stronger benefit of dataset scaling relative to model scaling may be explained by the much larger reductions in test loss achieved via data scaling, as seen by comparing Figures 3(c,g) with Figures 3(b,f), and quantified in Figure 2. The weak sensitivity to model size in iEEG may reflect the smaller effective sample size available in iEEG pretraining.
Practical Takeaway.
@@
These findings have direct relevances for resource allocation in EFM. Because the domain is data-constrained, the optimal compute strategy differs from that of language or vision: smaller models trained for more epochs outperform larger models trained briefly under fixed compute budgets. The IsoLoss and IsoCompute structure in Figure 3(p, left) indicates that large models are compute-inefficient at realistic epoch counts. Building on this observation, Figure 3(p, right) defines a clear “compute-optimal frontier” that specifies the most efficient combinations of model size and training duration. Prior EFMs tend to fall outside this frontier, often emphasizing increased model size over additional training repetitions (Appendix G). Our scaling framework provides a principled insight for selecting compute-optimal points before launching expensive pretraining. However, when the goal is to achieve the highest possible performance for a given model size, training beyond the compute-optimal frontier is still beneficial. This was the case for our Small iEEG model, and also applied to CBraMod (Wang et al., 2024c) and BIOT (Yang et al., 2023), when trained substantially past their compute-optimal points.
@@ 4.2 Decoding Performance
As shown in Figure 3(q)-(t), DIVER-1 achieves state-of-the-art decoding performance across iEEG and EEG benchmarks. On the iEEG downstream dataset, our 13M-parameter model surpasses nearly all prior approaches—including BrainBERT (43M) (Wang et al., 2023), PopT (63M)(Chau et al., 2025), and Brant (506M) (Zhang et al., 2023)—across almost all 15 tasks in Neuroprobe binaray-label tasks and by a large margin in MAYO. For Neuroprobe Multi-label tasks, our model still outperforms the linear STFT laplacian which surpasses other models. On EEG downstream tasks, DIVER-1 also establishes new SOTA results, outperforming CBraMod (Wang et al., 2024c) and LaBraM (Jiang et al., 2024) on FACED emotion recognition (Chen et al., 2023), PhysioNet-MI motor imagery (Goldberger et al., 2000), and MentalArithmetic workload decoding (Zyma et al., 2019). Full task-level values for all benchmark results are provided in Appendix Section D.1. We also evaluated DIVER-1 on additional EEG downstream tasks, where we identified a methodological (reproduction) problem. To ensure fair comparison and facilitate replicability, we conducted a controlled investigation by standardizing the finetuning method across models and comparing performance over seven different methods. We find that under controlled finetuning settings, DIVER-1 also achieves comparable or superior performance to the other baseline models (Appendix Table 25). Linear probing analysis further demonstrates that DIVER learns high-quality representations (Appendix Section E.5). Detailed experimental procedures and results are provided in Appendix Section E.
The strong performances over NeuroProbe are particularly notable given the pretraining–finetuning distribution shift under which DIVER-1 was trained. Unlike PopT or BrainBERT, DIVER-1 was pretrained on different datasets than those used for finetuning, on adult data rather than pediatrics, and using both ECoG and SEEG modalities rather than exclusively SEEG data. Moreover, DIVER-1 is much smaller (13M) than the other models (BrainBERT : 43M, PopT : 63M). This highlights the robustness of our representations across dataset shifts and recording modalities.
@@ 4.3 Ablation Studies
To disentangle architectural contributions to performance from the pretraining data-scale, we evaluate all models pretrained on the same dataset (BrainTreebank; Table 1). Under this controlled setting, DIVER still outperforms BrainBERT and PopT on downstream tasks, indicating that the gains are not solely attributable to larger pretraining corpora.
Table 1: Controlled comparison using BrainTreebank (BTB). Models are pretrained on the same BTB corpus to isolate architectural effects. Results are mean AUROC $\pm$ SEM across subjects, trials, and folds (DIVER: 32 pretraining epochs).
Model speech onset
BrainBERT (frozen)* $0.611\pm 0.022$ $0.757\pm 0.027$
PopT* $0.677\pm 0.044$ $0.689\pm 0.050$
[…]
* results from Neuroprobe (Zahorodnii et al., 2025)
Architecture ablations (Appendix Tables 18, 19) indicate most components are contributing to the final performance, including any-variate attention, RoPE, STCPE, and the multi-domain reconstruction objectives. The main exception was the iEEG absolute channel-position embedding, whose ablation improved performance. We hypothesize this reflects a pretraining–downstream distribution shift: pretraining uses adult population data, whereas Neuroprobe consists only of child/adolescent subjects, so age-related differences in brain size/geometry and electrode localization may miscalibrate absolute 3D positional embeddings, reducing their transferability and decreasing performance.
Figure 3: Scaling laws and downstream performance of DIVER-1. (a-h) Scaling law validation: DIVER-1 follows data-constrained scaling laws across four dimensions for iEEG (a-d) and EEG (e-h) modalities. Loss decreases predictably with increased (a,e) compute (training FLOPs), (b,f) dataset size (number of tokens), (c,g) model size (parameters), and (d,h) training epochs, with strong log-log fits. iEEG experiments (a-d) used 100% of the dataset, while EEG experiments used 20% of the dataset for (e,h) and 100% for (f,g). (q-t) Downstream performance: Performance for iEEG (i-l) and EEG (m-o) across increasing (i) number of subjects while keeping dataset size identical, (j,m) dataset size (k,n) model sizes and (l,o) epochs. (p) Compute-Optimal Frontier (IsoLoss analysis): Comparison between empirical isolation loss contours and predicted isolation loss contours, with model configurations plotted to show the relationship between training epochs and model parameters under fixed compute budgets. (q) Neuroprobe benchmark results Comprehensive performance (AUROC) comparison across multiple neural decoding tasks, with $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ achieving state-of-the-art or competitive results on most tasks. $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ with $d_{\text{model}}=256$ and patch size 0.1s was pretrained on iEEG dataset for 32 epochs, past the compute optimal frontier for best performance. Performance with linear probing (red) and full finetuning (black) are shown. (r, s) iEEG downstream performance (r) Neuroprobe multi-label classification results using $\textsc{DIVER}_{\text{Small}/\text{I}/0.1s}$. (s) MAYO(seizure detection task) results using $\textsc{DIVER}_{\text{Small}/\text{I}/1s}$ (t). EEG downstream performance: DIVER-1 showed competitive performance compared to other EEG foundation models (CBraMod and LaBraM-base) on the FACED, PhysioNet-MI, and MentalArithmetic datasets. Results shown are obtained using full finetuning. The DIVER model refers to $\textsc{DIVER}_{\text{Small}/\text{IE}/1s}$ with $d_{\text{model}}=512$ and patch size 1s pretrained on iEEG and EEG datasets for 16 epochs. Other baseline results are replicated using their official code. Performance values for CBraMod and LaBraM are reported from their original publications.